Find the ama bounded by awives x=34 and 9-1 = 64
Answer #1

Teh-zi£ – 26 ) = Favo (zle=h7-6), Bipa pl bizom, f = rumon che = sumon 11 - ife-49-6 = utbiur () Is A = {- liler 3y?< x 59-64